A satellite is in elliptical orbit with a period of 9.23 x 104 s about a planet of mass 5.50 x 1024 kg. At aphelion, at radius 5.4 x 107 m, the satellite’s angular speed is 7.657 x 10-5 rad/s.What is its angular speed at perihelion?
Answer #1

Aus! As the Satellite is experiencing no Jorgle, ence, angular mpmentum remains constant. Lpetiteon = Lapletion Ipwop = lawa.= 6.67X10(m2/kg 52) x 5.50 X 1024xgx (9.23X104)2 472 = • (19.245 X 105+24+8) (79.245 x 162) 3a = 4.23 2107 m , Radius of aphelion (Rp) = 2a-Ra . = 2*4.23 x 10 m -(504x10 m) = 3.06 x 107 m Plugging the value in up = ( k